OpenStudy (anonymous):

find all numbers for which the rational expression is undefined (m^3-6m)/(m^2-81)

jhonyy9 (jhonyy9):

- firstly when (m*2 -81) will be zero that so m=+/- 9 - second when m*3-6m=0 --- m(m*2-6)=0 --- m=0 m*2-6=0 --- m*2=6 --- m=+/- radical 6